OpenStudy (goformit100):

As they say, beggars can't be choosers, in fact begger take what they can get. A begger on the street can make one cigarette out of every 6 cigarette butts he finds. After one whole day of searching and checking public ashtrays the begger finds a total of 72 cigarette butts. How many cigarettes can he make and smoke from the butts he found?

OpenStudy (goformit100):

@Koikkara

OpenStudy (anonymous):

12

OpenStudy (dmezzullo):

Ok so for every 6 butts he finds he gets one cigar right?

OpenStudy (anonymous):

72 ÷ 6 = 12

OpenStudy (goformit100):

yes

OpenStudy (anonymous):

76/2=12

OpenStudy (goformit100):

Thankyou @dmezzullo @LoveYou*69

OpenStudy (anonymous):

srry 76 /6 =12